OpenStudy (anonymous):

In right triangle ABC shown to the right AC=3x and BC=5x. If the perimeter of triangle ABC is 84, then x=?

OpenStudy (anonymous):

find AB by pythagoras theorem... \[\Large (5x)^2-(3x)^2=AB^2\] \[\Large 16x^2=AB^2\] \[\Large 4x=AB\] now find perimeter...

OpenStudy (anonymous):

Thank you!